the square root of x-4 - the square root of x+8 =2
Answer by edjones(8007)

sqrt(x-4)-sqrt(x+8)=2
sqrt(x-4)-2=sqrt(x+8)
x-4-4sqrt(x-4)+4=x+8 square both sides.
-4sqrt(x-4)=8
(-4sqrt(x-4))^2=8^2
16(x-4)=64
16x-64=64
16x=128
x=8
.
Check:
sqrt(8-4)-sqrt(8+8)=2
2-4=2 False
There is no answer to this equation.
.
Ed